### v1

@@ 6.2 Proof of results in Section 3
Then for every $N\in\mathbb{N}$ and $\lambda>0$ it holds that for every $t\geq 0$, $\mathbb{P}$-a.s.,
$\big|\alpha_{t}^{*,x;N}-\pi^{*,x;N,\lambda}_{t}\big|$ $\leq\bigg|{\bf 1}_{\{D_{t}^{x;N}<0\}}-{\bf 1}_{\{\overline{D}_{t}^{x;N,\lambda}<0\}}\bigg|+\bigg|{\bf 1}_{\{\overline{D}_{t}^{x;N,\lambda}<0\}}-\frac{1}{1+e^{\frac{N}{\lambda}\overline{D}_{t}^{x;N,\lambda}}}\bigg|$
(6.23) $={\bf 1}_{\{D_{t}^{x;N}<0\leq\overline{D}_{t}^{x;N,\lambda}\}}+\frac{1}{1+e^{{N}{\lambda}^{-1}|\overline{D}_{t}^{x;N,\lambda}|}},$
where the last equality holds as $D_{t}^{x;N}\leq\overline{D}_{t}^{x;N,\lambda}$, $\mathbb{P}$-a.s., for all $t\geq 0$ (see (3.10)).
By Theorem 3.5, for any $N\in\mathbb{N}$ $\|Y^{x;N}-\overline{Y}^{x;N,\lambda}\|_{\mathbb{S}^{2}}=\|D^{x;N}-\overline{D}^{x;N,\lambda}\|_{\mathbb{S}^{2}}\to 0$ as $\lambda\downarrow 0$. This implies that for any $N\in\mathbb{N}$, $|D_{t}^{x;N}-\overline{D}_{t}^{x;N,\lambda}|\to 0$ $\mathbb{P}\otimes dt$-a.e. as $\lambda\downarrow 0$.
Comining this with the a priori estimates given in (6.23), we have for any $N\in\mathbb{N}$
$\big|\alpha_{t}^{*,x;N}-\pi^{*,x;N,\lambda}_{t}\big|\to 0\quad\mbox{$\mathbb{P}\otimes dt$-a.e., as $\lambda\downarrow 0$.}$
Furthermore, since $\big|\alpha_{t}^{*,x;N}-\pi^{*,x;N,\lambda}_{t}\big|\leq 2$, $\mathbb{P}\otimes dt$-a.e., for all $N\in\mathbb{N}$ and $\lambda>0$ (noting that $(\alpha^{*,x;N})_{N\in\mathbb{N}}\subseteq{\mathcal{}A}$ and $(\pi^{*,x;N,\lambda})_{N\in\mathbb{N},\lambda>0}\subseteq\Pi$), the dominated convergence theorem guarantees that the convergence in (3.12) holds for all $N\in\mathbb{N}$.
@@ 6.3 Proof of results in Section 4
Proof 6.6 (Proof of Theorem 4.1).
@@
$\begin{aligned} n_{t}:=&\frac{1}{\Delta Y_{t}}\Big(\overline{F}_{t}^{{n+1}}(\overline{Y}_{t}^{n+1},\overline{Z}_{t}^{n+1})-\overline{F}_{t}^{{n+1}}(\overline{Y}_{t}^{{n}},\overline{Z}_{t}^{{n+1}})\Big){\bf 1}_{\{\Delta Y_{t}\neq 0\}},\\ m_{t,i}:=&\frac{1}{\Delta Z_{t,i}}\Big(\overline{F}^{{n+1}}_{t}(\overline{Y}_{t}^{n+1},(\overline{Z}_{t,1}^{n},\dots,\overline{Z}_{t,i-1}^{n},\overline{Z}_{t,i}^{n+1},\dots,\overline{Z}_{t,d}^{n+1})^{\top})\\ &\quad\quad\quad-\overline{F}^{{n+1}}_{t}(\overline{Y}_{t}^{n+1},(\overline{Z}_{t,1}^{n},\dots,\overline{Z}_{t,i}^{n},\overline{Z}_{t,i+1}^{n+1},\dots,\overline{Z}_{t,d}^{n+1})^{\top})\Big){\bf 1}_{\{\Delta Z_{t,i}\neq 0\}}.\end{aligned}$
@@
$\Delta Y_{t}=\int_{t}^{T}\left(n_{s}\Delta Y_{s}+m_{s}^{\top}\Delta Z_{s}+\phi_{s}\right)ds-\int_{t}^{T}\Delta Z_{s}dB_{s}.$
@@
Therefore, by letting $\Gamma_{t}:=\exp(\int_{0}^{t}m_{s}dB_{s}+\int_{0}^{t}(-n_{s}-\frac{1}{2}|m_{s}|^{2})ds)$ for $t\in[0,T]$, applying Itô’s formula into $(\Gamma_{t}\Delta Y_{t})_{t\in[0,T]}$ and taking the conditional expectation $\mathbb{E}_{t}[\cdot]$,
@@
Set $C_{1}:=C_{\beta}+\kappa+N>0$. By the a priori estimate in [70, Theorem 4.2.3], there exists some $C_{2}>0$ (that depends on $C_{1},T,d$ but not on $n,\lambda$), such that [^7] For any $t\in[0,T]$ and $Y\in\mathbb{S}^{2}(\mathbb{R})$, denote by $\|Y\|_{\mathbb{S}^{2}_{t}}^{2}:=\mathbb{E}[\sup_{s\in[t,T]}|Y_{s}|^{2}]$. In analogy, for any $Z\in\mathbb{L}^{2}(\mathbb{R}^{d})$, denote by $\|Z\|^{2}_{\mathbb{L}^{2}_{t}}:=\mathbb{E}[\int_{t}^{T}|Z_{s}|^{2}ds]$.
@@
$\|\Delta^{n+1}\overline{Y}\|_{\mathbb{S}^{2}}^{2}+\|\Delta^{n+1}\overline{Z}\|_{\mathbb{L}^{2}}^{2}\leq C_{3}\int_{t}^{T}\|\Delta^{n}\overline{Y}\|_{\mathbb{S}^{2}_{t_{n}}}^{2}dt_{n}$
@@
$\qquad\leq\cdots$
$\qquad\leq({C}_{3})^{n}\int_{0}^{T}\int_{t_{n}}^{T}\cdots\int_{t_{2}}^{T}\|\Delta^{1}\overline{Y}\|_{\mathbb{S}^{2}_{t_{1}}}^{2}dt_{1}\cdots dt_{n-1}\;dt_{n}$
@@
The monotonicity of $\pi^{n+1}_{t}$ as $n\uparrow\infty$ is obvious from the logistic functional form on $\overline{Y}^{x;N,\lambda,\pi^{n}}$, which completes the proof.
@@
Proof 6.8.
We start with proving (i). According to [19, Theorem 8.9], it suffices to show that the generator $\check{F}_{\cdot}^{N,\lambda,\check{\pi}}(\cdot,\cdot,\cdot)$ given in (4.2) satisfies the condition (M1b) given in [19] (noting that $\check{X}^{t,x}$ given in (6.25) satisfies (M1f) therein; see Remark 2.4). Note that $\beta_{t}$ and $\check{\pi}_{t}(x)$ are uniformly bounded (see Setting 4), and $g$ is uniformly Lipschitz w.r.t. $z$ (see Definition 2.1). Therefore, $\check{F}_{\cdot}^{N,\lambda,\check{\pi}}(\cdot,\cdot,\cdot)$ is uniformly Lipschitz w.r.t. $(y,z)$ with the corresponding Lipschitz constant depending only on $C_{\beta},\lambda,N$ (not on $t,x$). Moreover, for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$,
@@
We now prove (ii). As $r(x),R(x),\check{\pi}_{t}(x)$ are continuous w.r.t $x$ for all $t\in[0,T]$, the condition ($\mathrm{M1b^{c}}$) given in [19] holds true. Therefore, an application of [19, Theorem 8.12] ensures that $v^{N,\lambda,\check{\pi}}(t,x):=\check{Y}_{t}^{t,x;N,\lambda,\check{\pi}}$ for $(t,x)\in[0,T]\times\mathbb{R}^{d}$ is a viscosity solution of the PDE given in the statement (ii); see (6.25). Moreover, using the flow property of $\{\check{X}_{s}^{t,x};t\leq s\leq T,x\in\mathbb{R}^{d}\}$ and the uniqueness of the solution of (6.25), we have for $t\leq s\leq T$, $\mathbb{P}\otimes ds$-a.e., $v^{N,\lambda,\check{\pi}}(s,\check{X}_{s}^{t,x})=\check{Y}_{s}^{s,\check{X}_{s}^{t,x};N,\lambda,\check{\pi}}=\check{Y}_{s}^{t,x;N,\lambda,\check{\pi}},$ that is, the property in (6.26) holds. Lastly, the regularity of $v^{N,\lambda,\check{\pi}}$ follows from the argument in the proof of [19, Theorem 8.12], which employs the $L^{p}$-estimation techniques in the proof of [50, Lemma 2.1 and Corollary 2.10].
Proof 6.9 (Proof of Corollary 4.3).

### v2

@@ 6.2 Proof of results in Section 3
Moreover, we let $\tau_{t}^{n}:=\inf\{s\in[t,T]:\mathcal{Y}_{s}^{n}\leq L_{s}\}\wedge T$, $\tau_{t}:=\inf\{s\in[t,T]:\mathcal{Y}_{s}\leq L_{s}\}\wedge T,$ $t\in[0,T]$. Then it holds for every $t\in[0,T]$ that as $n\uparrow\infty$, $\tau_{t}^{n}\uparrow\tau_{t}$ -$\mathbb{P}\text{-a.s.}$.
Proof 6.6.
By the condition (i), it holds that $\tau_{t}^{n}\leq\tau_{t}^{n+1}\leq\tau_{t}$ -$\mathbb{P}\text{-a.s.}$, for any $n\in\mathbb{N}$. Moreover, since the filtration $\mathbb{F}$ is right-continuous (see Section 1.2), there exists an $\mathbb{F}$-stopping time $\bar{\tau}_{t}$ such that $\bar{\tau}_{t}=\lim_{n\to\infty}\tau_{t}^{n}\leq\tau_{t},$ -$\mathbb{P}\text{-a.s.}$.
It suffices to show that $\bar{\tau}_{t}\geq\tau_{t}$ $\mathbb{P}$-a.s. To this end, from the conditions (i) and (ii), we consider a set $\Omega_{0}\subseteq\Omega$ satisfying that $\mathbb{P}(\Omega_{0})=1$ and every $\omega\in\Omega_{0}$ satisfies
$\mathcal{Y}_{s}^{n}(\omega)\leq\mathcal{Y}_{s}^{n+1}(\omega)\leq\mathcal{Y}_{s}(\omega)\quad\mbox{for all $s\in[t,T]$ and $n\in\mathbb{N}$},\quad\mbox{and $\Delta^{n}(\omega)\downarrow 0$ as $n\uparrow\infty$.}$
Now, fix $\omega\in\Omega_{0}$. If $\tau_{t}(\omega)=t$, then, by definition of $\tau_{t}$, we have $\mathcal{Y}_{t}(\omega)\leq L_{t}(\omega)$. This implies $\mathcal{Y}_{t}^{n}(\omega)\leq L_{t}(\omega)$ for all $n\in\mathbb{N}$, and therefore $\tau_{t}^{n}(\omega)=t$ for all $n\in\mathbb{N}$. As a consequence, we have $\bar{\tau}_{t}(\omega)=t=\tau_{t}(\omega).$
Next, suppose that $\tau_{t}(\omega)>t$. Let $\varepsilon\in(0,\tau_{t}(\omega)-t)$. By definition of $\tau_{t}(\omega)$,
$\mathcal{Y}_{s}(\omega)>L_{s}(\omega)\quad\text{for all }s\in[t,\tau_{t}(\omega)-\varepsilon].$
Since the map $t,\tau_{t}(\omega)-\varepsilon]\ni s\mapsto\mathcal{Y}_{s}(\omega)-L_{s}(\omega)$ is continuous, we can define
$m_{\omega,\varepsilon}:=\min_{s\in[t,\tau_{t}(\omega)-\varepsilon]}(\mathcal{Y}_{s}(\omega)-L_{s}(\omega))>0.$
We let $N_{\omega,\varepsilon}\in\mathbb{N}$ so that $\Delta^{n}(\omega)<m_{\omega,\varepsilon}$ for all $n\geq N_{\omega,\varepsilon}$. Then for every $n\geq N_{\omega,\varepsilon}$,
$\mathcal{Y}_{s}^{n}(\omega)\geq\mathcal{Y}_{s}(\omega)-\Delta^{n}(\omega)>L_{s}(\omega)\quad\mbox{for all $s\in[t,\tau_{t}(\omega)-\varepsilon]$.}$
This implies that $\tau_{t}^{n}(\omega)\geq\tau_{t}(\omega)-\varepsilon$ for all $n\geq N_{\omega,\varepsilon}$. By letting $n\to\infty$, we have $\bar{\tau}_{t}(\omega)\geq\tau_{t}(\omega)-\varepsilon.$ Since $\varepsilon>0$ is arbitrary, we conclude that $\bar{\tau}_{t}(\omega)\geq\tau_{t}(\omega).$
Since this inequality holds for any $\omega\in\Omega_{0}$ and we have $\mathbb{P}(\Omega_{0})=1$ (by definition of $\Omega_{0}$), it follows that $\bar{\tau}_{t}\geq\tau_{t}$ $\mathbb{P}$-a.s.. This completes the proof.
Proof 6.7 (Proof of Theorem 3.6).
Step 1. We start with showing that for any $\lambda>0$, $N\in\mathbb{N}$ and $t\in[0,T]$, $\overline{Y}_{t}^{x;N,\lambda}-\lambda(T-t)\log 2\leq Y_{t}^{x}$ $\mathbb{P}$-a.s..
For notational simplicity, let $(\Delta\overline{Y}^{x;N,\lambda},\Delta\overline{Z}^{x;N,\lambda}):=(\overline{Y}^{x;N,\lambda}-Y^{x},\overline{Z}^{x;N,\lambda}-Z^{x})$ and $(\Delta Y^{x;N},\Delta Z^{x;N}):=({Y}^{x;N}-Y^{x},{Z}^{x;N}-Z^{x})$.
Using (2.4) and (3.8) as well as (3.7), we have
$\Delta\overline{Y}_{t}^{x;N,\lambda}$ $=\int_{t}^{T}\big(-\beta_{s}\Delta\overline{Y}_{s}^{x;N,\lambda}+g(s,\overline{Z}_{s}^{x;N,\lambda})-g(s,Z_{s}^{x})+G_{s}^{x;N,\lambda}(\overline{Y}_{s}^{x;N,\lambda})\big)ds$
(6.19) $\quad-\int_{t}^{T}\Delta\overline{Z}_{s}^{x;N,\lambda}dB_{s}-(K_{T}^{x}-K_{t}^{x}).$
Moreover, using (2.4) and (2.7), we have
$-(K_{T}^{x}-K_{t}^{x})$ $=\int_{t}^{T}\big(\beta_{s}\Delta Y_{s}^{x;N}-g(s,Z_{s}^{x;N})+g(s,Z_{s}^{x})-N(R(X_{s}^{x})-Y_{s}^{x;N})^{+}\big)ds$
(6.20) $\quad+\int_{t}^{T}\Delta Z_{s}^{N}dB_{s}+\Delta Y_{t}^{x;N}.$
Substituting (6.20) into (6.19) gives
$\Delta\overline{Y}_{t}^{x;N,\lambda}$ $=\Delta Y_{t}^{x;N}-\int_{t}^{T}\beta_{s}(\overline{Y}_{s}^{x;N,\lambda}-Y_{s}^{x;N})ds$
(6.21) $\quad+\int_{t}^{T}\big(G_{s}^{x;N,\lambda}(\overline{Y}_{s}^{x;N,\lambda})-N(R(X_{s}^{x})-Y_{s}^{x;N})^{+}\big)ds$
$\quad+\int_{t}^{T}\big(g(s,\overline{Z}_{s}^{x;N,\lambda})-g(s,Z_{s}^{x;N})\big)ds-\int_{t}^{T}(\overline{Z}_{s}^{x;N,\lambda}-Z_{s}^{x;N})dB_{s}.$
From Remark 2.9 (iii) and (3.10) in Theorem 3.5, we have that for every $t\in[0,T]$
(6.22) $\Delta Y_{t}^{x;N}\leq 0\quad\mbox{and}\quad\overline{Y}_{t}^{x;N,\lambda}-Y_{t}^{x;N}\geq 0\quad\mbox{$\mathbb{P}$-a.s.}.$
Moreover, since $\underline{G}_{s}^{x;N,\lambda}(y):=G_{s}^{x;N,\lambda}(y)-N(R(X_{s}^{x})-y)^{+}\leq\lambda\log 2$ for every $y\in\mathbb{R}$ and $s\in[t,T]$, and $\overline{Y}_{t}^{x;N,\lambda}\geq Y_{t}^{x;N}$ $\mathbb{P}$-a.s., we have that for all $s\in[t,T]$,
(6.23) $G_{s}^{x;N,\lambda}(\overline{Y}_{s}^{x;N,\lambda})-N(R(X_{s}^{x})-Y_{s}^{x;N})^{+}\leq\underline{G}_{s}^{x;N,\lambda}(\overline{Y}_{s}^{x;N,\lambda})\leq\lambda\log 2\quad\mbox{$\mathbb{P}$-a.s..}$
From (6.22) and (6.23) as well as the $\kappa$-Lipschitz continuity of $g$ (see Definition 2.1 (ii)), it follows that $\Delta\overline{Y}_{t}^{x;N,\lambda}$ given in (6.21) satisfies, $\mathbb{P}$-a.s.,
$\Delta\overline{Y}_{t}^{x;N,\lambda}-\lambda(T-t)\log 2\leq\int_{t}^{T}\kappa|\overline{Z}_{s}^{x;N,\lambda}-Z_{s}^{x;N}|ds-\int_{t}^{T}(\overline{Z}_{s}^{x;N,\lambda}-Z_{s}^{x;N})dB_{s}.$
Since $Z^{x;N},\overline{Z}^{x;N,\lambda}\in\mathbb{L}^{2}(\mathbb{R}^{d})$ (see Remark 2.9 (i) and Theorem 3.4), the right-hand side of the above inequality is a ${\mathcal{}E}^{-\kappa}$-martingale. Hence, its conditional ${\mathcal{}E}^{-\kappa}$-expectation at $t$ equals zero $\mathbb{P}$-a.s.. This proves the claim.
Step 2: Proof of (i). We show that $R(X_{\cdot}^{x})$, $Y^{x}$, and $(Y_{t}^{x;N_{n}}-\lambda_{n}(T-t)\log 2)_{t\in[0,T]}$, $n\in\mathbb{N}$, satisfy all the conditions of Lemma 6.5, so that for every $t\in[0,T]$,
(6.24) $\underline{\tau}_{t}^{n}:=\inf\{s\in[t,T]:Y_{s}^{x;N_{n}}-\lambda_{n}(T-s)\log 2\leq R(X_{s}^{x})\}\wedge T$
goes to $\tau_{t}^{*,x}$ $\mathbb{P}$-a.s., as $n\to\infty$ (see (2.5)).
First, we note that all those processes are $\mathbb{F}$-progressively measurable, continuous processes because $X^{x}$, $Y^{x}$, and $Y^{x;N_{n}}$, $n\in\mathbb{N}$, are $\mathbb{F}$-progressively measurable and continuous (see (2.1), (2.4), (2.7)), and $R$ is continuous (see Assumption 2 (i)).
Since $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$ as $n\to\infty$, by Remark 2.9 (iii), we have that for every $n\in\mathbb{N}$ and $t\in[0,T]$, $Y^{x;N_{n}}_{t}-\lambda_{n}(T-t)\log 2\leq Y^{x;N_{n+1}}_{t}-\lambda_{n+1}(T-t)\log 2\leq Y_{t}^{x}$ $\mathbb{P}$-a.s.. Therefore, the condition (i) in Lemma 6.5 holds.
Next, we set $\Delta^{x,n}:=\sup_{t\in[0,T]}(Y_{t}^{x}-(Y^{x;N_{n}}_{t}-\lambda_{n}(T-t)\log 2))\geq 0$, which is non-increasing $\mathbb{P}$-a.s.. Define $\Delta^{x,\infty}:=\lim_{n\to\infty}\Delta^{x,n}$. Then by Fatou’s lemma,
$\mathbb{E}\big[|\Delta^{x,\infty}|^{2}\big]\leq\liminf_{n\to\infty}\mathbb{E}\big[|\Delta^{x,n}|^{2}\big]\leq\liminf_{n\to\infty}\big(\|Y^{x}-Y^{N_{n}}\|_{\mathbb{S}^{2}}^{2}+\lambda_{n}T\log 2\big)=0,$
where the last equality follows from Remark 2.9. Thus, the condition (ii) holds.
Consequently, all the conditions in Lemma 6.5 are satisfied, we have by the lemma that $\underline{\tau}_{t}^{n}\uparrow\tau_{t}^{*,x}$ $\mathbb{P}$-a.s. as $n\to\infty$.
Last, we have by Step 1 and (3.10) that for any $n\in\mathbb{N}$ and $t\in[0,T]$,
$Y^{x;N_{n}}_{t}-\lambda_{n}(T-t)\log 2\leq\overline{Y}^{x;N_{n},\lambda_{n}}_{t}-\lambda_{n}(T-t)\log 2\leq Y_{t}^{x}\quad\mbox{$\mathbb{P}$-a.s..}$
This implies that for any $n\in\mathbb{N}$ and $t\in[0,T]$, $\underline{\tau}_{t}^{n}\leq\tau_{t}^{x;N_{n},\lambda_{n}}\leq\tau_{t}^{*,x}$ (see (6.24)). Thus, the convergence, $\underline{\tau}_{t}^{n}\uparrow\tau_{t}^{*,x}$ $\mathbb{P}$-a.s., ensures the statement (i) to hold.
Step 3: Proof of (ii). Since $\tau_{t}^{x;N_{n},\lambda_{n}}\to\tau_{t}^{*,x}$ $\mathbb{P}$-a.s. (by Step 2) and $R$ is continuous, by the properties of the integrand $\operatorname{I}_{t}^{x;\tau}$ for any $\tau\in{\mathcal{}T}_{t}$ and $t\in[0,T]$ given in Remark 2.5, we have that for any $t\in[0,T]$, $\|\operatorname{I}_{t}^{x;\tau_{t}^{x;N_{n},\lambda_{n}}}-\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}\|_{L^{2}}\to 0$ as $n\to\infty$.
Moreover, since $\operatorname{I}_{t}^{x;\tau}\in L^{2}({\mathcal{}F}_{T})$ for any $\tau\in{\mathcal{}T}_{t}$ and $t\in[0,T]$, by [pardoux1990adapted, Theorem 3.1] we can define $(\widetilde{Y}^{t,x;\tau},\widetilde{Z}^{t,x;\tau})\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ as a unique solution of
(6.25) $\widetilde{Y}_{s}^{t,x;\tau}=\operatorname{I}_{t}^{x;\tau}+\int_{s}^{T}g(u,\widetilde{Z}_{u}^{t,x;\tau})du-\int_{s}^{T}\widetilde{Z}_{u}^{t,x;\tau}dB_{u},\quad s\in[0,T].$
Clearly, it holds that for all $t\in[0,T]$, $\mathbb{P}$-a.s.,
$\widetilde{Y}_{t}^{t,x;\tau_{t}^{x;N_{n},\lambda_{n}}}={\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{x;N_{n},\lambda_{n}}}]\quad\mbox{and}\quad\widetilde{Y}_{t}^{t,x;\tau_{t}^{x;*}}={\mathcal{}E}^{g}_{t}[\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}]=V_{t}^{x}.$
Since $g$ is $\kappa$-Lipschitz, the parameters of the BSDEs (6.25) satisfy the conditions given in [el1997backward, Section 5] (with exponent $2$) under any $\tau\in{\mathcal{}T}_{t}$ and $t\in[0,T]$, we are able to apply [el1997backward, Proposition 5.1] to have the following a priori estimate: there exists $C>0$ (depends on $\kappa$ and $T$, but not on $n$) such that for any $n\in\mathbb{N}$ and $t\in[0,T]$,
$\big\|{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{x;N_{n},\lambda_{n}}}]-V_{t}^{x}\big\|_{L^{2}}\leq\|\widetilde{Y}^{t,x;\tau_{t}^{x;N_{n},\lambda_{n}}}-\widetilde{Y}^{t,x;\tau_{t}^{*,x}}\|_{\mathbb{S}^{2}}\leq C\|\operatorname{I}_{t}^{x;\tau_{t}^{x;N_{n},\lambda_{n}}}-\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}\|_{L^{2}}.$
Since the last term vanishes as $n\to\infty$, this completes the proof.
@@ 6.3 Proof of results in Section 4
Proof 6.8 (Proof of Theorem 4.1).
@@
$\begin{aligned} n_{t}:=&\frac{1}{\Delta Y_{t}}\Big(\overline{F}_{t}^{{n+1}}(\overline{Y}_{t}^{n+1},\overline{Z}_{t}^{n+1})-\overline{F}_{t}^{{n+1}}(\overline{Y}_{t}^{{n}},\overline{Z}_{t}^{{n+1}})\Big){\bf 1}_{\{\Delta Y_{t}\neq 0\}},\\ m_{t,i}:=&\frac{1}{\Delta Z_{t,i}}\Big(\overline{F}^{{n+1}}_{t}(\overline{Y}_{t}^{n},(\overline{Z}_{t,1}^{n},\dots,\overline{Z}_{t,i-1}^{n},\overline{Z}_{t,i}^{n+1},\overline{Z}_{t,i+1}^{n+1},\dots,\overline{Z}_{t,d}^{n+1})^{\top})\\ &\quad\quad\quad-\overline{F}^{{n+1}}_{t}(\overline{Y}_{t}^{n},(\overline{Z}_{t,1}^{n},\dots,\overline{Z}_{t,i-1}^{n},\overline{Z}_{t,i}^{n},\overline{Z}_{t,i+1}^{n+1},\dots,\overline{Z}_{t,d}^{n+1})^{\top})\Big){\bf 1}_{\{\Delta Z_{t,i}\neq 0\}}.\end{aligned}$
@@
$\Delta Y_{t}=\int_{t}^{T}(n_{s}\Delta Y_{s}+m_{s}^{\top}\Delta Z_{s}+\phi_{s})ds-\int_{t}^{T}\Delta Z_{s}dB_{s}.$
@@
Therefore, by letting $\Gamma_{t}:=\exp(\int_{0}^{t}m_{s}dB_{s}+\int_{0}^{t}(n_{s}-\frac{1}{2}|m_{s}|^{2})ds)$ for $t\in[0,T]$, applying Itô’s formula into $(\Gamma_{t}\Delta Y_{t})_{t\in[0,T]}$ and taking the conditional expectation $\mathbb{E}_{t}[\cdot]$,
@@
Set $C_{1}:=C_{\beta}+\kappa+N>0$. By the a priori estimate in [zhangjianfeng17, Theorem 4.2.3], there exists some $C_{2}>0$ (that depends on $C_{1},T,d$ but not on $n,\lambda$), such that [^5] For any $t\in[0,T]$ and $Y\in\mathbb{S}^{2}(\mathbb{R})$, denote by $\|Y\|_{\mathbb{S}^{2}_{t}}^{2}:=\mathbb{E}[\sup_{s\in[t,T]}|Y_{s}|^{2}]$. In analogy, for any $Z\in\mathbb{L}^{2}(\mathbb{R}^{d})$, denote by $\|Z\|^{2}_{\mathbb{L}^{2}_{t}}:=\mathbb{E}[\int_{t}^{T}|Z_{s}|^{2}ds]$.
@@
$\|\Delta^{n+1}\overline{Y}\|_{\mathbb{S}^{2}}^{2}+\|\Delta^{n+1}\overline{Z}\|_{\mathbb{L}^{2}}^{2}\leq C_{3}\int_{0}^{T}\|\Delta^{n}\overline{Y}\|_{\mathbb{S}^{2}_{t_{n}}}^{2}dt_{n}$
@@
$\qquad\leq\cdots\leq({C}_{3})^{n}\int_{0}^{T}\int_{t_{n}}^{T}\cdots\int_{t_{2}}^{T}\|\Delta^{1}\overline{Y}\|_{\mathbb{S}^{2}_{t_{1}}}^{2}dt_{1}\cdots dt_{n-1}\;dt_{n}$
@@
The monotonicity of $\pi^{n+1}_{t}$ as $n\uparrow\infty$ is obvious from the logistic function form on $\overline{Y}^{x;N,\lambda,\pi^{n}}$, which completes the proof.
@@
Proof 6.10.
We start with proving (i). According to [KHM, Theorem 8.9], it suffices to show that the generator $\check{F}_{\cdot}^{N,\lambda,\check{\pi}}(\cdot,\cdot,\cdot)$ given in (4.2) satisfies the condition (M1b) given in [KHM] (noting that $\check{X}^{t,x}$ given in (6.27) satisfies (M1f) therein; see Remark 2.4). Note that $\beta_{t}$ and $\check{\pi}_{t}(x)$ are uniformly bounded (see Setting 4), and $g$ is uniformly Lipschitz w.r.t. $z$ (see Definition 2.1). Therefore, $\check{F}_{\cdot}^{N,\lambda,\check{\pi}}(\cdot,\cdot,\cdot)$ is uniformly Lipschitz w.r.t. $(y,z)$ with the corresponding Lipschitz constant depending only on $C_{\beta},\lambda,N$ (not on $t,x$). Moreover, for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$,
@@
We now prove (ii). As $r(x),R(x),\check{\pi}_{t}(x)$ are continuous w.r.t $x$ for all $t\in[0,T]$, the condition ($\mathrm{M1b^{c}}$) given in [KHM] holds true. Therefore, an application of [KHM, Theorem 8.12] ensures that $v^{N,\lambda,\check{\pi}}(t,x):=\check{Y}_{t}^{t,x;N,\lambda,\check{\pi}}$ for $(t,x)\in[0,T]\times\mathbb{R}^{d}$ is a viscosity solution of the PDE given in the statement (ii); see (6.27). Moreover, using the flow property of $\{\check{X}_{s}^{t,x};t\leq s\leq T,x\in\mathbb{R}^{d}\}$ and the uniqueness of the solution of (6.27), we have for $t\leq s\leq T$, $\mathbb{P}\otimes ds$-a.e., $v^{N,\lambda,\check{\pi}}(s,\check{X}_{s}^{t,x})=\check{Y}_{s}^{s,\check{X}_{s}^{t,x};N,\lambda,\check{\pi}}=\check{Y}_{s}^{t,x;N,\lambda,\check{\pi}},$ that is, the property in (6.28) holds. Lastly, the regularity of $v^{N,\lambda,\check{\pi}}$ follows from the argument in the proof of [KHM, Theorem 8.12], which employs the $L^{p}$-estimation techniques in the proof of [pardoux2005backward, Lemma 2.1 and Corollary 2.10].
Proof 6.11 (Proof of Corollary 4.5).